# Multimodal Fish Feeding Intensity Assessment in Aquaculture

### v2

@@ Abstract
Feeding is one of the most important costs in aquaculture. However, current feeding machines usually operate with fixed thresholds or human experiences, lacking the ability to automatically adjust to fish feeding intensity. FFIA can evaluate the intensity changes in fish appetite during the feeding process and optimize the control strategies of the feeding machine to avoid inadequate feeding or overfeeding, thereby reducing the feeding cost and improving the well-being of fish in industrial aquaculture. The existing methods have mainly exploited single-modality data, and have a high sensitivity to input noise. Using video and audio offers improved chances to address the challenges brought by various environments. However, compared with processing data from single modalities, using multiple modalities simultaneously often involves increased computational resources, including memory, processing power, and storage. This can impact system performance and scalability. To address these issues, we focus on the efficient unified model, which is capable of processing both multimodal and single-modal input. Our proposed model achieved SOTA performance in FFIA with high computational efficiency.
@@ I Introduction
Digital aquaculture plays an important role in enhancing operational efficiency, sustainability, and productivity in aquaculture [1, 2]. As the foundation of digital aquaculture technology, Fish Feeding Intensity Assessment (FFIA) is the task of evaluating the fish feeding intensity during the feeding process. FFIA offers significant benefits for real-world aquaculture by facilitating adjustments to feed dispensing, minimizing waste, and ultimately improving productivity [3].
Traditional methods of FFIA have relied on the experience of human observers and subjective interpretation [4], which can be time-consuming and labour-intensive. With the recent advances in deep learning, data-driven approaches [5] have become the emerging methods for FFIA, such as images based fish feeding intensity classification (e.g., “None”, “Weak”, “Medium” and “Strong[6]), using Convolutional Neural Networks (CNNs), and video-based FFIA methods [7, 8]. In comparison, video-based methods offer advantages in capturing spatial and temporal information of the feeding process, which offers richer context for capturing fish feeding behaviour, as compared with the methods based on single images [9]. Nevertheless, processing video data can involve more computational resources, which may make it less practical for on-device applications that play a crucial role in aquaculture [10]. Furthermore, vision-based methods are prone to changes in lighting conditions and noise resulting from water surface reflections [11].
As an alternative to vision-based methods, acoustic measurements can also be used for FFIA. In comparison, acoustic measurements are unaffected by illumination changes and occlusions, suitable for round-the-clock 24 hours monitoring, and are often more cost-effective to process [12, 13]. However, acoustics-based methods can be limited in capturing the physical characteristics related to fish behaviour, and are prone to underwater acoustic noises, thereby degrading the FFIA performance.
@@ II Related work
Vision-based FFIA methods have initially considered single images, and employed traditional techniques, such as background segmentation and target tracking [18]. However, these methods suffered from high computational loads and were prone to environmental factors like water surface fluctuations [11]. More recent works have considered recognition tasks including fish classification [19], counting [20], and tracking [21] by leveraging CNN models. For FFIA, models such as MSIF-MobileNetV3 [22] and EfficientNet-B2 [23] have been successfully applied.
@@
The fusion of audio and visual modalities has the potential to address the limitations posed by each modality. This has been demonstrated in various fields, such as audio-visual event localization [34], audio-visual speech recognition [35], audio-visual emotion recognition [36], and animal behaviour analysis [37]. However, to the best of our knowledge, no previous work has explored the fusion of audio-visual information for FFIA, which is the focus of our work in this paper.
@@ III-B AV Fusion via Multi-Head Cross-Attention
In contrast to Self-Attention, Multi-Head Cross-Attention (MHCA) allows for the integration of both audio and visual information, leveraging the strengths of each modality. The MHCA operation can be written as:
@@ IV-D Unified Model with Knowledge Distillation
Due to the limited number of frames used for both audio and visual inputs, the performance of the U-FFIA model can be compromised with the use of the single model architecture. To address this limitation, we employ knowledge distillation (KD) to enhance the performance of our audio and visual models.
Knowledge distillation [16] is a model compression technique in which a large-scale model (teacher) is compressed into a smaller model (student) while preserving similar performance. The key idea is to use the logits output by the teacher model (in our case, the pre-trained audio model CNN10 [15] and visual model S3D [49]) as the supervision information to guide the learning of the smaller student model. The teacher model uses all the audio and visual frames, while the student model uses a limited number of frames. During training, the teacher model is frozen, and the focus is solely on training the student model using the knowledge distillation loss. We use the following loss for the student model training:
@@ VI-A1 Audio Data Processing
We use mel spectrogram as acoustic features, which has been widely used for audio classification [15]. The original audio sampling rate is $256$ kHz, which can lead to computational and storage burdens. We down-sample the audio signals to $64$ kHz to reduce the computational complexity. Then we calculate the Short-Time Fourier Transform (STFT) with a Hanning window of $2048$ samples and a hop size of $1024$ samples, and finally, we apply the mel filter banks with $128$ bins. Therefore, for a $2$-second audio signal, we have a mel spectrogram with the shape of $128$$\times$ $128$. We use the SimPF method [14] which is a simple non-parametric pooling operation, to reduce the redundant information within the mel-spectrogram, leading to a shape of $64$$\times$ $128$ mel spectrogram. This streamlined mel spectrogram representation enhances the efficacy of audio classification tasks. To evaluate the noise robustness of the model, we add bubble and pump noise to the entire audio signal with different Signal-to-Noise Ratio (SNR) levels, ranging from -$10$ dB to $20$ dB. This allows us to assess the model’s performance under varying noise conditions.
@@ VII-A The Benchmark Results of Audio-Based FFIA
TABLE III: The Performance of different AV fusion methods on AV-FFIA dataset and AV fusion under noisy environment.
@@
The results also indicated that pre-trained models PANNs [15] and AST [52] outperformed the models trained from scratch, i.e. MobileNetV1 [63], ResNet18 [65] highlighting the benefits of transfer learning for the FFIA task. Transformer models i.e. Vanilla [43], originally designed for sequential data in natural language processing (NLP) tasks, heavily rely on Self-Attention mechanisms. While these mechanisms excel in modelling long-range dependencies, they might struggle to discern fine-grained local patterns evident in audio spectrograms, leading to a suboptimal performance on the AV-FFIA dataset compared to CNN-based models.
@@ VII-C The Benchmark Results of Audio-Visual Based FFIA
To illustrate this point, we can consider the fish feeding scenario, where the fish may quickly consume all the current feed, resulting in no sound being produced. However, the video still shows the fish highly aggregated and in a state of hunger. In such cases, the cross-attention method allows the information from the audio modality to complement that from the visual modality, enabling the model to make more accurate predictions. On the other hand, the MBT [45] method achieved a lower accuracy of only $0.724$ compared to the cross-attention method. This performance gap could be attributed to attention bottlenecks that attend to each modality sequentially, leading to a failure to capture the complex relationships among modalities. Consequently, there may be information loss and a lack of direct interaction between modalities, which can limit the model’s ability to effectively fusing the audio and visual information.
@@ VII-F Experiments in Real Aquaculture Environment
To show the performance of our proposed method in real-world scenarios, we conducted experiments in a commercial aquaculture facility (as shown in Fig. 5). To this end, we collected a total of 12,700 audio-visual samples from an aquaculture environment at the Guangzhou Aquatic Products Promotion Station, China. The real scene contains more environmental noise, water surface reflection, foam, and other challenges, as compared with the controlled environment.
The real-world dataset focuses on Tilapia fish, a commonly farmed species in aquaculture. We used two identical tanks, each measuring $4$ meters in length, $2$ meters in width, and $3$ meters in depth, to house the Tilapia fish during the data collection process. The real-world dataset follows the same specifications and equipment setting as the controlled dataset, with each sample consisting of a $2$-second audio clip and its corresponding video clip. The samples were manually annotated by experienced aquaculture technicians to provide ground truth labels for model evaluation. We randomly split the real-world dataset into training ($70$%), validation ($10$%), and testing ($20$%) sets.
We evaluated our proposed method on the real-world dataset and compared its performance with the results obtained in the controlled experiments. We found that the video modality achieved an accuracy of $0.827$, while the audio modality achieved an accuracy of $0.813$. Our proposed U-FFIA model still can achieve an accuracy of $0.886$, which is still higher than the state-of-the-art models (Cross-Attention [67], Self-Attention [43], and MBT [45] ). We observed that in the noisy environment, the video performance decreased significantly due to the darkness and overlap of the spume. However, the audio performance did not decrease as much because the pump and water flow operate at stable frequencies, which are quite different from the frequencies of fish slapping water or ingesting feed. This difference in frequencies allows the audio modality to maintain its performance even in the presence of background noise.
@@ IX ACKNOWLEDGMENT
This work was supported by the Research and demonstration of digital cage integrated monitoring system based on underwater robot [China grant 2022YFE0107100], Digital Fishery Cross-Innovative Talent Training Program of the China Scholarship Council (DF-Project) and a Research Scholarship from the China Scholarship Council. Ethics approval for this study was obtained from the Welfare and Ethical Committee of China Agricultural University (Ref: AW30901202-5-1). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author-accepted manuscript version arising.

### v3

@@ Abstract
Feeding is one of the most important costs in aquaculture. However, current feeding machines usually operate with fixed thresholds or human experiences, lacking the ability to automatically adjust to fish feeding intensity. FFIA can evaluate the intensity changes in fish appetite during the feeding process and optimize the control strategies of the feeding machine to avoid inadequate feeding or overfeeding, thereby reducing the feeding cost and improving the well-being of fish in industrial aquaculture. The existing methods have mainly exploited single-modality data, and have a high sensitivity to input noise. Using video and audio offers improved chances to address the challenges brought by various environments. However, compared with processing data from single modalities, using multiple modalities simultaneously often involves increased computational resources, including memory, processing power, and storage. This can impact system performance and scalability. To address these issues, we focus on the efficient unified model, which is capable of processing both multimodal and single-modal input. Our proposed model achieved state-of-the-art (SOTA) performance in FFIA with high computational efficiency.
@@ I Introduction
Digital aquaculture plays an important role in enhancing operational efficiency, sustainability, and productivity in aquaculture [1]. As the foundation of digital aquaculture technology, fish feeding intensity assessment (FFIA) is the task of evaluating the fish feeding intensity during the feeding process [2]. FFIA offers significant benefits for real-world aquaculture by facilitating adjustments to feed dispensing, minimizing waste, and ultimately improving productivity [3].
Traditional FFIA methods have relied on human observers’ experience and subjective interpretation [4], which can be time-consuming and labour-intensive. With the recent advances in deep learning, data-driven approaches [5] have emerged as promising methods for FFIA, such as image-based fish feeding intensity classification (e.g., “None”, “Weak”, “Medium” and “Strong[6]) using convolutional neural networks (CNNs) [7], and video-based FFIA methods [8, 9]. Video-based methods offer advantages in capturing spatial and temporal information of the feeding process, providing a richer context for fish feeding behaviour analysis compared to single-image methods [10]. However, processing video data often requires greater computational resources, which may make it less practical for on-device applications that play a crucial role in aquaculture [11]. Furthermore, vision-based methods are generally susceptible to variations in lighting conditions and noise from water surface reflections [12], affecting their reliability in real-world aquaculture settings.
As an alternative to vision-based methods, acoustic measurements can also be used for FFIA. In comparison, acoustic measurements are unaffected by illumination changes and occlusions, suitable for round-the-clock 24 hours monitoring, and are often more cost-effective to process [13, 14]. However, acoustics-based methods can be limited in capturing the physical characteristics related to fish behaviour, and are prone to underwater acoustic noises, potentially compromising FFIA performance.
@@ II Related work
Vision-based FFIA methods have initially considered single images, and employed traditional techniques, such as background segmentation and target tracking [19]. However, these methods suffered from high computational loads and were prone to environmental factors like water surface fluctuations [12]. More recent works have considered recognition tasks including fish classification [20], counting [21], tracking [22, 23] and trawl fishing [24] by leveraging CNN or Transformer models. For FFIA, models such as MSIF-MobileNetV3 [25] and EfficientNet-B2 [26] have been successfully applied.
@@
The fusion of audio and visual modalities has the potential to address the limitations posed by each modality. This has been demonstrated in various fields, such as audio-visual event localization [37], audio-visual speech recognition [38], audio-visual emotion recognition [39], and animal behaviour analysis [40, 41]. However, to the best of our knowledge, no previous work has explored the fusion of audio-visual information for FFIA, which is the focus of our work in this paper.
@@ III-B AV Fusion via Multi-Head Cross-Attention
In contrast to self-attention, multi-head cross-attention (MHCA) [49, 50] allows for the integration of both audio and visual information, leveraging the strengths of each modality. The MHCA operation can be written as:
@@ IV-D Unified Model with Knowledge Distillation
Due to the limited number of frames used for both audio and visual inputs, the performance of the U-FFIA model can be compromised with the use of the single model architecture. To address this limitation, we employ knowledge distillation (KD) [55] to enhance the performance of our audio and visual models.
Knowledge distillation [17] is a model compression technique in which a large-scale model (teacher) is compressed into a smaller model (student) while preserving similar performance. The key idea is to use the logits output by the teacher model (in our case, the pre-trained audio model CNN10 [16] and visual model S3D [56]) as the supervision information to guide the learning of the smaller student model. The teacher model uses all the audio and visual frames, while the student model uses a limited number of frames. During training, the teacher model is frozen, and the focus is solely on training the student model using the knowledge distillation loss.
We use the following loss for the student model training:
@@ VI-A1 Audio Data Processing
We use mel spectrogram as acoustic features, which has been widely used for audio classification [16]. The original audio sampling rate is $256$ kHz, which can lead to computational and storage burdens. We down-sample the audio signals to $64$ kHz to reduce the computational complexity. Then, we calculate the short-time Fourier transform (STFT) with a Hanning window of $2048$ samples and a hop size of $1024$ samples, and finally, we apply the mel filter banks with $128$ bins. Therefore, for a $2$-second audio signal, we have a mel spectrogram with the shape of $128$$\times$ $128$. We use the SimPF method [15] which is a simple non-parametric pooling operation, to reduce the redundant information within the mel-spectrogram, leading to a shape of $64$$\times$ $128$ mel spectrogram. This streamlined mel spectrogram representation enhances the efficacy of audio classification tasks. To evaluate the noise robustness of the model, we add bubble and pump noise to the entire audio signal with different SNR levels, ranging from -$10$ dB to $20$ dB. This allows us to assess the model’s performance under varying noise conditions.
@@ VII-A The Benchmark Results of Audio-Based FFIA
TABLE III: The Performance of different AV fusion methods on the AV-FFIA dataset and AV fusion under noisy environment.
@@
The results also indicated that pre-trained models PANNs [16] and AST [59] outperformed the models trained from scratch, i.e. MobileNetV1 [70], ResNet18 [72] highlighting the benefits of transfer learning for the FFIA task. The transformer models, such as [47], which were originally designed for sequential data in natural language processing (NLP) tasks, are built on the self-attention mechanisms. While these mechanisms excel in modelling long-range dependencies, they might struggle to discern fine-grained local patterns evident in audio spectrograms, leading to suboptimal performance on the AV-FFIA dataset compared to CNN-based models.
@@ VII-C The Benchmark Results of Audio-Visual Based FFIA
To illustrate this point, we can consider the fish feeding scenario, where the fish may quickly consume all the current feed, resulting in no sound being produced. However, the video still shows the fish highly aggregated and in a state of hunger. In such cases, the cross-attention method allows the information from the audio modality to complement that from the visual modality, enabling the model to make more accurate predictions. On the other hand, the MBT [51] method achieved a lower accuracy of only $0.724$ compared to the cross-attention method. This performance gap could be attributed to attention bottlenecks that attend to each modality sequentially, leading to a failure to capture the complex relationships among modalities. Consequently, there may be information loss and a lack of direct interaction between modalities, which can limit the model’s ability to effectively fuse the audio and visual information.
@@ VII-F Experiments in Real Aquaculture Environment
To demonstrate the real-world applicability of our method, we conducted experiments with a commercial aquaculture facility at the Guangzhou Aquatic Products Promotion Station, China (as shown in Fig. 5). We collected 12,700 audio-visual samples of Tilapia fish, a common aquaculture species, from two identical tanks ($4$m x $2$m x $3$m). This real-world environment presented challenges such as environmental noise, water surface reflection, and foams, which are not present in controlled settings. The dataset maintained the same specifications as our controlled dataset, with $2$-second audio-visual clips annotated by experienced technicians. We randomly split the real-world dataset into training ($70$%), validation ($10$%), and testing ($20$%) sets.
To adapt our model for this real data, we employed a two-stage training approach. First, we pretrained our model on the AV-FFIA dataset and then fine-tuned it on real-world data. This approach allowed us to leverage the large-scale AV-FFIA dataset while adapting the model to the specific conditions of the real aquaculture environment. Our fine-tuned U-FFIA model achieved $0.892$ accuracy on the real-world dataset, outperforming the SOTA models (e.g., cross-attention [74], self-attention [47], and MBT [51]). We observed that on the real dataset, the video performance ($0.822$) decreased significantly due to the darkness and overlap of the foams. However, the audio performance ($0.816$) did not decrease as much because the pump and water flow operate at stable frequencies, which are quite different from the frequencies of fish slapping water or ingesting feed. This frequency difference allows the audio modality to maintain its performance even in the presence of background noise.
@@ IX ACKNOWLEDGMENT
This work was supported by the National Natural Science Foundation of China “Dynamic regulation mechanism of nitrogen in industrial aquaponics under the condition of asynchronous life cycle” [Grant no. 32373186], Digital Fishery Cross-Innovative Talent Training Program of the China Scholarship Council (DF-Project) and a Research Scholarship from the China Scholarship Council. Ethics approval for this study was obtained from the Welfare and Ethical Committee of China Agricultural University (Ref: AW30901202-5-1). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any author-accepted manuscript version arising.